OpenStudy (anonymous):

What is the solution to the equation 49x=sqrt 7^(2x+6)

OpenStudy (saifoo.khan):

\[\Large 49x = \sqrt{ 7^{(2x+6)}}\]Like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Take square on both sides, you will get \[\large (48x)^2 = 7^{2x+6}\]

OpenStudy (saifoo.khan):

That's 49 not 48

OpenStudy (anonymous):

ohhh now i remember that lesson. Thanks :)

OpenStudy (saifoo.khan):

\[\Large (49x)^2 = 7^{2(x+3)}\] \[\Large (7^2x)^2 = 7^{2(x+3)}\]

OpenStudy (saifoo.khan):

\[\Large 49x = 7^{(x+3)}\]